OpenStudy (anonymous):

Answer this for a fan and a medal! Sylvia teaches music to 63 students who are under 10 years of age. The students under 10 years of age make up 35% of all her music students. How many students does sylvia teach in total?

OpenStudy (anonymous):

63 + 35% =?

OpenStudy (anonymous):

what he typed...

OpenStudy (anonymous):

Thanks @victorhernandez

OpenStudy (anonymous):

np :P

OpenStudy (anonymous):

but not, 63 students are 35%, so 100% will be....

OpenStudy (anonymous):

you want to do 63/35%

OpenStudy (anonymous):

so you do 63 + 35% then divide that answer by 100% or is it multiply? :3

OpenStudy (anonymous):

no... 63/0.35=180 total number of students is 180

OpenStudy (anonymous):

hmmm seems interesting can be confusing .-. errm

OpenStudy (anonymous):

If only i can medal two ppl... FIGHT TO THE DEATH THE ONE WITHE ANSWR WINS DA MEDAL!!!!

OpenStudy (anonymous):

but 35% are only under 10

OpenStudy (anonymous):

dude...just do it like this 63 ------- 0.35 x ------- 1 => x=63*1/0.35=180

OpenStudy (anonymous):

mhm :3

OpenStudy (anonymous):

winner!

OpenStudy (anonymous):

oooh ok I shouldn't done it like u XD

OpenStudy (anonymous):

should of*

OpenStudy (anonymous):

lol thanks anyway @victorhernandez

OpenStudy (anonymous):

I learned something new XD

OpenStudy (anonymous):

lol and u got a fan